OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

2x + 3y = 1,200 3y = -2x + 1,200 y = x + 400 Slope: m = y-intercept: b = 400 MF Question i had this on a test about a week ago

OpenStudy (anonymous):

thank you I just got so confused :)

OpenStudy (anonymous):

i did to but my teacher helped me work through it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

do you think you could help me with another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok I got that answer but now how would I describe how I would graph this line using the slope-intercept method. Be sure to write in complete sentences

OpenStudy (anonymous):

To graph this line, I would start by identifying the point where the line crosses the y-axis. Since the y-intercept is 400, the line crosses at (0, 400). The slope is , so the run is 2 and the rise is -3. I would draw a point at (0, 400), and then move 2 units to the right and 3 units down to plot the next point. I would draw a line through the two points.

OpenStudy (anonymous):

okay I think I got it now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank so much:)

OpenStudy (anonymous):

anytime